# Leveraging Multimodal LLM for Inspirational User Interface Search

### v2

@@ 1. Introduction
User interface (UI) design plays a central role in creating intuitive and engaging digital experiences. A vital part of the UI design process is the inspirational search, where designers explore existing UI designs to inform and inspire their own work (Herring et al., 2009; Siangliulue et al., 2015). By leveraging a wide array of reference materials, designers can craft innovative, user-centered interfaces that build upon established design patterns and best practices (Chan et al., 2015; Gonçalves et al., 2014; Koch et al., 2019; Bonnardel, 1999). During an inspirational search process for UI design, designers often focus on functionalities, problem domains, and visual styles (Lu et al., 2022). These aspects offer valuable insights into the design choices of UIs and assist designers in comprehending how to fulfill user needs and expectations effectively.
However, current UI retrieval methods are limited in their ability to capture the nuanced semantic elements essential to an effective, inspirational design process. Prior research (Huang et al., 2019; Wei et al., 2024) have focused on pixel-level retrieval methods, emphasizing visual similarities. While these approaches are effective for identifying surface-level design patterns, they often overlook in-depth semantic aspects such as user interaction flow, target audience, and the emotional tone of an application. Some computational methods (Liu et al., 2018; Li et al., 2021; Bunian et al., 2021) attempt to address these limitations by incorporating metadata like view hierarchies and text labels to enable more comprehensive browsing of UI semantics. However, the effectiveness of these methods is constrained by the quality and availability of metadata, which can vary widely across different UI datasets and design tools, limiting their scalability and applicability. Likewise, platforms like Pinterest, Behance, and Dribbble are popular due to their vast collections of design examples and are frequently used for visual inspiration. However, these platforms primarily offer image-based content with limited ability to filter by deeper semantic attributes, making it challenging for designers to find contextually relevant UI references (Wu et al., 2021). UI curation platforms like Mobbin, which categorize designs by app information, screen category, and UI elements, offer some improvements but rely heavily on manual curation and human annotation, providing limited scalability.
@@
To explore the potential of this approach, we began our research by investigating the following key research questions:
  • RQ1: What semantics are most important to UI designers during an inspirational search?
  • RQ2: How accurately can a multimodal large language model (MLLM) extract semantic elements from UI images?
  • RQ3: What are the advantages of a semantic-based UI search system over a traditional pixel or keyword-based retrieval?
To answer these questions, we first conducted a formative study with designers to identify the most crucial semantic facets of inspirational search. Based on these findings, we developed a robust methodology for extracting and structuring semantic information using MLLM. We then conducted a multi-stage evaluation process to assess the effectiveness of our semantic extraction pipeline. This evaluation included quantitative assessments using established UI datasets to measure the accuracy of app category classification and screen type identification, UI element prediction, and screen role understanding. We also performed qualitative evaluations with designers to validate the quality of extracted semantics and to assess their impact on the overall search experience. Furthermore, we compared our approach with existing pixel-based methods to demonstrate the advantages of our MLLM-based semantic search system across multiple aspects.
The main contributions of this paper are threefold:
@@
  • We propose a novel MLLM-based approach for extracting rich semantic information directly from UI images without relying on additional metadata.
  • We demonstrate the effectiveness of our approach through extensive evaluations, showing significant improvements over existing UI search methods in relevance, diversity, reliability, usefulness, and serendipity.
@@ 2.1. Inspirational Search in UI Design Process
The user interface design process is a complex and iterative endeavor that involves multiple stages, from ideation to implementation (Rogers et al., 2023). This complexity drives designers to search for external sources of inspiration that can inform their design decisions and stimulate creativity (Chan et al., 2015; Gonçalves et al., 2014). Herring et al. (Herring et al., 2009) highlighted the importance of examples as a source of inspiration in the design process, and multiple studies (Bonnardel, 1999; Siangliulue et al., 2015; Lee et al., 2010) noted their role in enhancing designers’ creativity and lead to more diverse and innovative solutions.
@@ 2.2. Computational UI Retrieval
Several works have focused on learning latent semantic embeddings for UI components and screens, enabling more effective retrieval and clustering of similar UIs. Deka et al. (Liu et al., 2018) introduced a method for learning semantically meaningful embeddings of UI components, allowing for efficient search and analysis of UI design patterns. Bunian et al. (Bunian et al., 2021) introduced VINS, a UI design search system that utilizes wireframes or screenshots to query and retrieve relevant UI examples. Huang et al. (Huang et al., 2019) proposed Swire, a similar system that combines sketch-based retrieval with UI component recognition to enable more flexible and intuitive search experiences. Methods like Screen2vec (Li et al., 2021) generate embeddings that integrate UI components, layouts, and texts for robust retrieval. These approaches enabled retrieval of semantically similar screens by training latent representation of semantics. However, they are utilizing not only UI images but also metadata, such as view hierarchies, which are often noisy (Li et al., 2022) and inherently limit their scalability. Recognizing this challenge, recent machine learning approaches have shown another direction. Park and Kim et al. (Park et al., 2023) demonstrated that large-scale foundation models like CLIP (Radford et al., 2021), trained without relying on additional metadata, can achieve superior retrieval performance compared to traditional approaches. As a stretching of this approach, GUIClip (Wei et al., 2024), a CLIP model fine-tuned with UI datasets, also showed better retrieval performance than prior methods.
@@ 2.3. Multimodal Learning and Language Models in UI Design
More recent studies have begun to explore the application of MLLMs on UI that go beyond the simple vision-language model. For example, Ferret-UI (You et al., 2025) combines a customized visual sampler with MLLMI structure, enabling UI-related tasks such as referring, grounding, and reasoning of user interfaces. Likewise, AppAgent (Zhang et al., 2023) leverages an MLLM to operate the user interface in a manner similar to human users, enabling autonomous task execution. While these approaches demonstrate the potential of MLLMs in UI understanding, they are not specifically tailored for UI design tasks. Furthermore, most of the above approaches require not only images but additional UI metadata (e.g., types of UI elements or coordinates) at the time of inference. This reliance on metadata makes it impractical to search for UI design inspiration, as the vast majority of inspiration materials lack structured UI metadata.
This study addresses the challenge of effectively capturing and leveraging the semantic elements crucial for designers during inspirational search and creative processes, using only screen images. We propose a novel MLLM-based approach for semantic extraction and retrieval, enabling designers to capture the nuanced intentions underlying their work. Our study represents a step forward by demonstrating the utility of MLLMs for extracting designer-oriented semantic elements vital for inspirational UI design search, using only screenshot images.
@@ 3. Formative Study
3.1. Participants and Procedure
To gain a comprehensive understanding of UI inspirational search practices, we recruited six participants, balancing experienced professionals and emerging designers. This sample size was chosen based on prior research, indicating that it effectively captures recurring themes and meaningful insights within a focused qualitative study (Guest et al., 2006; Morse, 2000). Balancing expertise levels ensured a broad range of perspectives, which enriches the depth and applicability of findings (HU and CHANG, 2017; Cross, 2004). This diversity allowed us to capture general needs from early-career designers to experienced practitioners, providing a more comprehensive understanding of design practices. The final participants included three expert UI designers and three design major students with UI design experiences, as shown in Table 1.
@@ 3.1. Participants and Procedure
All interviews were audio-recorded with participant consent and later transcribed for analysis. We employed a thematic analysis approach, iteratively coding the transcripts to identify recurring themes, challenges, and desired features. The authors conducted semantic coding of the interview data, organizing insights into key categories relevant to the design process. This process allowed us to distill key insights that informed our understanding of essential semantic elements for UI search and inspiration.
This study received approval from the Institutional Review Board (IRB). Participants were fully informed of the study’s purpose and procedures, and their consent was obtained prior to participation. Participants received compensation equivalent to 15 USD for their time and contribution.
@@ 3.2.1. UI Design Process and Inspirational Search
Participants consistently highlighted the importance of the Double Diamond design process (Council, [n. d.]), which involves both problem definition and solution exploration. Early in the process, designers focus on understanding user needs, defining core features, and considering the app’s target audience (P1, 2, 4-6). One participant (P2), a designer with eight years of experience, remarked:
“In the discovery phase, we focus on understanding the user and defining the app’s features. Knowing the target audience shapes much of the design process.”
@@ 3.4. Deriving Design Principles
DP1: Considering both Functionality and Aesthetics
Existing inspiration platforms often prioritize aesthetics but overlook functional aspects of UI, while UI-specific platforms focus on functionality but lack support for aesthetics like mood and tone. Designers focus on functionality in the early stages and shift to aesthetics in later stages, but current systems fail to support both aspects effectively. This lack of integration of current tools hinders the iterative design process as designers struggle to revisit and refine their work holistically. Supporting both functionality and aesthetics in an integrated manner allows for richer inspiration at each stage, enabling designs that are both effective and visually compelling.
DP2: Flexible Query for Adjusting Priorities
Our formative study revealed that designers’ search priorities shift across different design stages, focusing more on functional elements in the early design stages and on aesthetic aspects in wireframing and visual refinement. This principle emphasizes the need for a search approach that adapts to these shifting priorities, enabling designers to focus on certain semantic elements while still exploring diverse options.
DP3: Adaptive and Exploratory Search
Designers emphasized the iterative nature of their search process and the difficulty of articulating precise keywords to find relevant references. This implies the need for iterative query refinement, including showing the semantics of search results, to overcome the limitations of keyword-based searches and facilitate adaptive and exploratory discovery during inspirational searches.
@@ 4.1. Semantic Extraction
4.1.1. Defining Categorical Semantics
4.1.2. Handling Layouts and Visual Design
For more complex design attributes, such as layout and mood, we curated semantic definitions from existing design literature and filtered them to ensure relevance to mobile interfaces. For the layout, we adopted UI layout guides from online material [^2] https://devsquad.com/blog/user-interface-layouts, curating a set of layouts specific to mobile UI and excluding those intended for desktop or other environments. We selected guideline articles because describing intricate layouts textually is inherently challenging (Lu et al., 2023), and these articles offer semantically meaningful textual interpretations of the layout. For mood extraction, we referenced Nielsen Norman Group’s guidelines on moodboard (Moran, 2016) to provide a robust set of mood keywords and filter those relevant to mobile design. For colors, we adopted a hybrid approach. Color is a critical visual design component, but current LLMs often struggle with precise color extraction (Wang et al., 2024). To address this issue in the context of color schemes, we leveraged LLMs to generate initial color scheme categories. This approach takes advantage of the extensive knowledge of the LLM of well-known color schemes, allowing it to generate schemes that align with the general design intent. For the color palette, taking inspiration from a previous study (Wang et al., 2023), which adopts HTML format for LLM’s UI understanding, we presented color information using HTML Color Names. ded structure to multi-faceted attributes such as mood and layout by assigning concise keys to each attribute where needed. For example, a screen’s mood can be represented by a key-value pair, where the key is the mood, and the value provides a brief explanation of its design implications (e.g., “playful: A simple and straightforward design facilitating focus on habits.”). Using keys helps organize attributes for faster lookup, while detailed descriptions support both embedding-based searches and the adaptive and exploratory search principle. Table 3 provides an overview of the structured output formats for all key UI semantics.
@@ 4.1.3. Prompting Strategy for MLLM
To achieve optimal results from the MLLM, we implemented a detailed prompting strategy following established prompt engineering principles that include five key components (Figure 2): Assistant Persona, Task Instruction, Feature List, Feature Definition and Instruction, and Response Form. Each of these was encapsulated in an XML format to enhance the efficiency of the prompt execution (Anthropic, 2023b). The Assistant Persona component defined the MLLM as a mobile application design expert specializing in interpreting mobile UI screenshots and identifying detailed characteristics. This approach is based on the role prompting technique  (Amatriain, 2024; Xu et al., 2023), which effectively frames the model’s responses within a specific context and expertise level. This persona was designed to draw insights from screenshots and provide expert-level analysis tailored to the specific context presented. The Task Instruction component outlined the assistant’s task as understanding the screen’s content, extracting text, identifying UI elements’ roles and positions, and recognizing relationships between them. We provided clear and specific instructions that are fundamental to effective prompt engineering, guiding the model’s focus and output (Amatriain, 2024; Google, 2024; Anthropic, 2023a). The Feature List component included a comprehensive set of semantic attributes as identified earlier. The Feature Definition and Instruction component provided detailed definitions and instructions for extraction for each feature. The model was instructed to respond with appropriate values within the predefined categories for categorical semantics. For example, in the case of UI elements, the model was tasked with identifying each element’s role, while for the next and previous screens, it was guided to align its output with the inferred screen role. For the color palette, the assistant was prompted to analyze the color distribution first and extract the primary and secondary colors. It was also instructed to output the colors of major UI components in HTML Color Name format, ensuring clarity and consistency in the response. This detailed approach to definitions and instructions aligns with the principle of providing comprehensive context to improve the model’s understanding and performance (Fagbohun et al., 2024; OpenAI, 2024; Anthropic, 2023a). Finally, the Response Form component specified the output format. We chose the output format as YAML, which was selected for its human readability and ability to reduce token counts up to 50% (Livshitz, 2023) by eliminating unnecessary characters like braces and reducing blank space indentation of JSON.

### v3

@@ 1. Introduction
User interface (UI) design plays a central role in creating intuitive and engaging digital experiences. A vital part of the UI design process is the inspirational search, where designers explore existing UI designs to inform and inspire their own work (Herring et al., 2009; Siangliulue et al., 2015). By leveraging a wide array of reference materials, designers can craft innovative, user-centered interfaces that build upon established design patterns and best practices (Chan et al., 2015; Gonçalves et al., 2014; Koch et al., 2019; Bonnardel, 1999). During inspirational search process for UI design, designers often focus on functionalities, problem domains, and visual styles (Lu et al., 2022). These aspects offer valuable insights into the design choices of UIs and assist designers in comprehending how to fulfill user needs and expectations effectively.
However, current UI retrieval methods are limited in their ability to capture the nuanced semantic elements essential to an effective, inspirational design process. Prior research (Huang et al., 2019; Wei et al., 2024) have focused on pixel-level retrieval methods, emphasizing visual similarities. While these approaches are effective for identifying surface-level design patterns, they often overlook in-depth semantic aspects such as user interaction flow, target audience, and the emotional tone of an application. Some computational methods (Liu et al., 2018; Li et al., 2021; Bunian et al., 2021) have attempted to address these limitations by incorporating metadata such as view hierarchies and text labels to enable more comprehensive browsing of UI semantics. However, the effectiveness of these methods is constrained by the quality and availability of metadata, which can vary widely across different UI datasets and design tools, limiting their scalability and applicability. Likewise, platforms like Pinterest, Behance, and Dribbble are popular due to their vast collections of design examples and are frequently used for visual inspiration. However, these platforms primarily offer image-based content with limited ability to filter by deeper semantic attributes, making it challenging for designers to find contextually relevant UI references (Wu et al., 2021). UI curation platforms like Mobbin, which categorize designs by app information, screen category, and UI elements, offer some improvements but rely heavily on manual curation and human annotation, providing limited scalability.
@@
To explore the potential of this approach, we begin our research by investigating the following key research questions:
  1. RQ1
    What UI semantics are most important for UI designers during their search for inspiration?
  2. RQ2
    How accurately can a multimodal large language model (MLLM) extract semantic elements from UI images?
  3. RQ3
    What are the advantages of semantic-based UI search compared to traditional pixel or keyword-based retrieval?
To answer these research questions, we first conducted a formative study with designers to identify the most crucial semantic facets of inspirational search. Based on these findings, we developed a novel methodology for extracting and structuring semantic information using MLLM. Utilizing these extracted semantics, we designed S&UI, a semantic-based UI retrieval system. To assess the effectiveness of our approach, we conducted a multi-stage evaluation process. This evaluation included computational assessments using established UI datasets to measure the accuracy of app and screen category classification, UI element prediction, and screen role understanding. We also performed human evaluations with designers, combining quantitative ratings and qualitative insights to validate the quality and usefulness of the extracted semantics. Furthermore, we compared our semantic-based UI search method to existing pixel and keyword-based retrieval methods. Through this comparative study, we demonstrated the advantages of our approach across multiple aspects, showcasing its effectiveness in inspirational UI search through quantitative ratings and qualitative insights from designer interviews.
Building on these efforts, we highlight three key contributions:
@@ 3. RQ3
  • We propose a novel MLLM-based approach for extracting rich semantics directly from UI screen images and assess its effectiveness through computational and human evaluations.
  • We introduce S&UI, a semantic-based UI retrieval system, and demonstrate its significant strengths over existing pixel and keyword-based UI search methods in multiple criteria.
@@ 2.1. Inspirational Search in UI Design Process
The user interface design process is a complex and iterative endeavor that involves multiple stages, from ideation to implementation (Rogers et al., 2023). This complexity drives designers to search for external sources of inspiration that can inform their design decisions and stimulate creativity (Chan et al., 2015; Gonçalves et al., 2014). Herring et al. (Herring et al., 2009) highlighted the importance of examples as a source of inspiration in the design process, and multiple studies (Bonnardel, 1999; Siangliulue et al., 2015; Lee et al., 2010) noted their role in enhancing designers’ creativity and leading to more diverse and innovative solutions.
@@ 2.2. Computational UI Retrieval
Several works have focused on learning latent semantic embeddings for UI components and screens, enabling more effective retrieval and clustering of similar UIs. Deka et al. (Liu et al., 2018) introduced a method for learning semantically meaningful embeddings of UI components, allowing for efficient search and analysis of UI design patterns. Bunian et al. (Bunian et al., 2021) introduced VINS, a UI design search system that utilizes wireframes or screenshots to query and retrieve relevant UI examples. Huang et al. (Huang et al., 2019) proposed Swire, a similar system that combines sketch-based retrieval with UI component recognition to enable more flexible and intuitive search experiences. Methods like Screen2Vec (Li et al., 2021) generate embeddings that integrate UI components, layouts, and texts for robust retrieval. These approaches enabled retrieval of semantically similar screens by training latent representation of semantics. However, they are utilizing not only UI images but also metadata, such as view hierarchies, which are often noisy (Li et al., 2022) and inherently limit their scalability. Recognizing this challenge, recent machine learning approaches have shown another direction. Park and Kim et al. (Park et al., 2023) demonstrated that large-scale foundation models like CLIP (Radford et al., 2021), trained without relying on additional UI-specific metadata, can achieve superior retrieval performance compared to traditional approaches. As a stretching of this approach, GUIClip (Wei et al., 2024), a CLIP model fine-tuned with UI datasets, also showed better retrieval performance than prior methods.
@@ 2.3. Multimodal Learning and Language Models in UI Design
More recent studies have begun to explore the application of MLLMs on UI that go beyond the simple vision-language model. For example, Ferret-UI (You et al., 2025) combines a customized visual sampler with MLLM, enabling UI-related tasks such as referring, grounding, and reasoning of user interfaces. Likewise, AppAgent (Zhang et al., 2023) leverages an MLLM to operate the user interface in a manner similar to human users, enabling autonomous task execution. While these approaches demonstrate the potential of MLLMs in UI understanding, they are not specifically tailored for UI design tasks. Furthermore, most of the above approaches require not only images but additional UI metadata (e.g., types of UI elements or coordinates) at the time of inference. This reliance on metadata makes it impractical to search for UI design inspiration, as the vast majority of inspiration materials lack structured UI metadata.
This study addresses the challenge of effectively capturing and leveraging the semantic elements crucial for designers during inspirational searches and creative processes, using only screen images. We propose a novel MLLM-based approach for semantic extraction and retrieval, enabling designers to capture the nuanced intentions underlying their work. Our study represents a step forward by demonstrating the utility of MLLMs for extracting designer-oriented semantic elements vital for inspirational UI design search, using only screenshot images.
@@ 3. Formative Study
3.1. Participants and Procedure
To gain a comprehensive understanding of UI inspirational search practices, we recruited six participants, balancing experienced professionals and emerging designers. This sample size was chosen based on prior research, which indicates that it effectively captures recurring themes and meaningful insights within a focused qualitative study (Guest et al., 2006; Morse, 2000). Balancing expertise levels ensured a broad range of perspectives, which enriches the depth and applicability of findings (HU and CHANG, 2017; Cross, 2004). This diversity allowed us to capture general needs from early-career designers to experienced practitioners, providing a more comprehensive understanding of design practices. The final participants included three expert UI designers and three design major students with UI design experiences, as shown in Table 1.
@@ 3.1. Participants and Procedure
All interviews were audio-recorded with participant consent and later transcribed for analysis. We employed thematic analysis, iteratively coding the transcripts to identify recurring themes, challenges, and desired features. The authors conducted semantic coding of the interview data, organizing insights into key categories relevant to the design process. This process allowed us to distill key insights that informed our understanding of essential semantic elements for UI search and inspiration.
This study received approval from the Institutional Review Board (IRB). Participants were fully informed of the study’s purpose and procedures, and their consent was obtained prior to participation. The participants received compensation equivalent to 15 USD for their time and contribution.
@@ 3.2.1. UI Design Process and Inspirational Search
Participants consistently highlighted the importance of the Double Diamond design process (Council, [n. d.]), which involves both the problem definition and the solution exploration phases. During the process, designers focus on understanding user needs, defining core features, and considering the target audience of the app (P1, 2, 4-6). One participant (P2), a designer with eight years of experience, remarked:
“In the discovery phase, we focus on understanding the user and defining the features of the app. Knowing the target audience shapes much of the design process.”
@@ 3.4. Deriving Design Principles
  1. DP1
    Considering both Functionality and Aesthetics. Existing inspiration platforms often prioritize aesthetics but overlook functional aspects of UI, while UI-specific platforms focus on functionality but lack support for aesthetics like mood and tone. Designers focus on functionality in the early stages and shift to aesthetics in later stages, but current systems fail to support both aspects effectively. This lack of integration of current tools hinders the iterative design process as designers struggle to revisit and refine their work holistically. Supporting both functionality and aesthetics in an integrated manner allows for richer inspiration at each stage, enabling designs that are both effective and visually compelling.
  2. DP2
    Flexible Query for Adjusting Priorities. Our formative study revealed that designers’ search priorities shift across different design stages, focusing more on functional elements in the early stages and on aesthetic aspects in wireframing and visual refinement. This principle emphasizes the need for a search approach that adapts to shifting priorities, allowing designers to emphasize certain semantics while maintaining a level of consistency in the overall search results.
  3. DP3
    Adaptive and Exploratory Search. Designers emphasized the iterative nature of their search process and the difficulty of articulating precise search keywords to find relevant references. This highlights the need for iterative query refinement, allowing designers to adjust their searches dynamically. Additionally, surfacing the semantics of search results can help overcome the limitations of keyword-based searches, facilitating a more adaptive and exploratory discovery process during inspirational searches.
@@ 4.1. Semantic Extraction
4.1.1. Defining Categorical Semantics
4.1.2. Handling Layouts and Visual Design
For more complex design attributes, such as layout and mood, we curated semantic definitions from existing design literature and filtered them to ensure relevance to mobile interfaces. For the layout, we adopted UI layout guides from online material [^2] https://devsquad.com/blog/user-interface-layouts, curating a set of layouts specific to mobile UI and excluding those intended for desktop or other environments. We selected guideline articles because describing intricate layouts textually is inherently challenging (Lu et al., 2023), and these articles offer semantically meaningful textual interpretations of the layout. For mood extraction, we referenced Nielsen Norman Group’s guidelines on moodboard (Moran, 2016) to provide a robust set of mood keywords and filter those relevant to mobile design. For colors, we adopted a hybrid approach. Color is a critical visual design component, but current LLMs often struggle with precise color extraction (Wang et al., 2024). To address this issue in the context of color schemes, we leveraged LLMs to generate initial color scheme categories. This approach takes advantage of the extensive knowledge of the LLM of well-known color schemes, allowing it to generate schemes that align with the general design intent. For the color palette, inspired by a previous study (Wang et al., 2023), which adopts HTML format for LLMs’ UI understanding, we presented color information using HTML color names. We added structure to multi-faceted attributes such as mood and layout by assigning concise keys to each attribute where needed. For example, a screen’s mood can be represented by a key-value pair, where the key is the mood, and the value provides a brief explanation of its design implications (e.g., “playful: A simple and straightforward design facilitating focus on habits.”). Using keys helps organize attributes for faster lookup, while detailed descriptions support both embedding-based searches and the adaptive and exploratory search principle. Table 3 provides an overview of the structured output formats for all key UI semantics.
@@ 4.1.3. Prompting Strategy for MLLM
To achieve optimal results from the MLLM, we implemented a detailed prompting strategy following established prompt engineering principles that include five key components (Figure 2): Assistant Persona, Task Instruction, Feature List, Feature Definition and Instruction, and Response Form. Each component was encapsulated in an XML format to enhance the efficiency of the prompt execution (Anthropic, 2023b). The Assistant Persona component defined the MLLM as a mobile application design expert specializing in interpreting mobile UI screenshots and identifying detailed characteristics. This approach is based on the role prompting technique  (Amatriain, 2024; Xu et al., 2023), which effectively frames the model’s responses within a specific context and expertise level. This persona was designed to draw insights from screenshots and provide expert-level analysis tailored to the specific context presented. The Task Instruction component outlined the assistant’s task as understanding the screen’s content, extracting text, identifying UI elements’ roles and positions, and recognizing relationships between them. We provided clear and specific instructions that are fundamental to effective prompt engineering, guiding the model’s focus and output (Amatriain, 2024; Google, 2024; Anthropic, 2023a). The Feature List component included a comprehensive set of semantic attributes as identified earlier. The Feature Definition and Instruction component provided detailed definitions and instructions for extraction for each feature. The model was instructed to respond with appropriate values within the predefined categories for categorical semantics. For example, in the case of UI elements, the model was tasked with identifying each element’s role, while for the next and previous screens, it was guided to align its output with the inferred screen role. For the color palette, the assistant was prompted to analyze the color distribution first and extract the primary and secondary colors. It was also instructed to output the colors of major UI components in HTML Color Name format, ensuring clarity and consistency in the response. This detailed approach to definitions and instructions aligns with the principle of providing comprehensive context to improve the model’s understanding and performance (Fagbohun et al., 2024; OpenAI, 2024; Anthropic, 2023a). Finally, the Response Form component specified the output format. We chose the output format as YAML, which was selected for its human readability and ability to reduce token counts up to 50% (Livshitz, 2023) by eliminating unnecessary characters like braces and reducing blank space indentation of JSON.